OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! A circle is growing so that each side is increasing at the rate of 5 cm/min. How fast is the area of the circle changing at the instant the radius is 20 cm? Include units in your answer.

OpenStudy (aryana_maria2323):

Can you please help? @across @Directrix @ganeshie8 @jim_thompson5910 @myininaya @satellite73 @zepdrix

OpenStudy (across):

A circle has no sides. Do you mean radius?

OpenStudy (aryana_maria2323):

Yes it wants to know the area of the circle changing at the instant the radius is 20cm

OpenStudy (across):

\[A=\pi r^2\\\frac{dA}{dt}=2\pi r\frac{dr}{dt}\] Solve for \(dA/dt\). You know what \(r\) and \(dr/dt\) are.

OpenStudy (aryana_maria2323):

I know r is 20 but what is dr/dt?

OpenStudy (across):

It's in the first sentence.

OpenStudy (aryana_maria2323):

Can you please show me steps? I am not exactly sure how to do it.

OpenStudy (across):

Just substitute\[r=20\\\frac{dr}{dt}=5\]

OpenStudy (aryana_maria2323):

so \[\frac{ d(20) }{ dt }=5\] Like this?

OpenStudy (across):

dude\[\frac{dA}{dt}=2\pi(20)(5)=200\pi\quad\frac{\text{cm.}^2}{\text{min.}}\]

OpenStudy (aryana_maria2323):

wait why is there an A where the r is?

OpenStudy (across):

You're looking for how fast area changes relative to radius. So you need an expression for the area of the circle in terms of its radius. That's easy. That is:\[A=\pi r^2\]. But I'm interested in how the area changes with respect to time, so I differentiate with respect to time (applying chain rule and all):\[\frac{dA}{dt}=2\pi r\frac{dr}{dt}.\]The problem tells me what \(r\) and \(dr/dt\) are, so it's just now a matter of plugging in numbers and that's it.